# Quantum-enabled optical large-baseline interferometry: applications, protocols and feasibility

### v1

@@ Appendix B Calculating the mean photon numbers of nearby stars and exoplanets
In this section, we include details of how we arrive at the mean photon numbers of nearby stars and exoplanets. Here the monochromatic absolute magnitude $m_{AB}$ is defined as the logarithm of a spectral flux density with the usual scaling of astronomical magnitudes The zero-point is approximately about 3631 Jy, where
@@
Denote $F_{\nu}$ as the spectral flux density
@@
Now, to calculate the total photon specific bandwidth, we assume the density is approximately constant over the bandwidth (this holds for narrow bandwidths):
@@
For the CHARA array, visible interferometry has bandwith between 10-50 nm.
@@
Table 7: Expected number of photons to collect at given wavelength, bandwidth for the host star. The wavelengths of interest are: 760nm is the absorption line for molecular oxygen; for Earth-like conditions the pressure-broadening linewidth is typically $\sim$ 2 pm at FWHM; 1.65 $\mu$m is a methane absorption line, and for earth-like conditions the FWHM ×$\times 10^{-4}\mu$m. Here we have assumed a collection area of 10 m2. Spectral demultiplexing doesn’t decrease $\epsilon$ due to the fact that that the effective time bins get larger. We assume a conservative $10^{-9}$ factor between the star and planet’s brightness. These numbers might differ depending on what assumption is made about the temperature of the star.

### v2

@@ Appendix A Calculating the mean photon numbers of nearby stars and exoplanets
In this section, we include details of how we arrive at the mean photon numbers of nearby stars and exoplanets. Here the monochromatic absolute magnitude $m_{AB}$ is defined as the logarithm of a spectral flux density with the usual scaling of astronomical magnitudes. The zero-point is approximately about 3631 Jy, where
@@
Denote $F_{\nu}$ as the spectral flux density.
@@
Now, to calculate the total photon-specific bandwidth, we assume the density is approximately constant over the bandwidth (this holds for narrow bandwidths):
@@
For the CHARA array, visible interferometry has a bandwidth between 10-50 nm.
@@
Table 5: Expected number of photons to collect at a given wavelength, bandwidth for the host star. The wavelengths of interest are: 760nm is the absorption line for molecular oxygen; for Earth-like conditions the pressure-broadening linewidth is typically $\sim$ 2 pm at FWHM; 1.65 $\mu$m is a methane absorption line, and for earth-like conditions the FWHM ×$\times 10^{-4}\mu$m. Here we have assumed a collection area of 10 m2. Spectral demultiplexing doesn’t decrease $\epsilon$ due to the fact that the effective time bins get larger. We assume a conservative $10^{-9}$ factor between the star and the planet’s brightness. These numbers might differ depending on what assumption is made about the temperature of the star.